The density of mercury is 13.6 g/ml. Calculate approximately the diameter of an atom of mercury assuming that each atom is occupying a cube of edge length equal to the diameter of the mercury atom.

Grade:10

1 Answers

Vikas TU
7 years ago
Volume of cube = d^3 where d is the diameter of atomHg.
Thus,
Volume = mass/density = 200/13.6 = 14.705 =  d^3
d = 2.4143